The complement of an angle measures 37°. What is the measure of the angle?
Answer by nyc_function(2733)

(90 - A) = complement of an angle

Let A = the angle

90 - A = 37

-A = 37 - 90

-A = -53

A = -53/-1

A = 53

The angle is 53 degrees.